The ratio of the measure of two suplementary angles is (4x-10)/(5x+1). Find the value of X.
Answer by jim_thompson5910(35256)

Let A and B be the two angles that are supplementary.

Let A = 4x-10
Let B = 5x+1


A+B = 180
(4x-10)+(5x+1) = 180
9x-9 = 180
9(x-9) = 180
x-9 = 180/9
x-9 = 20
x = 20+9
x = 29


The value of x is x = 29